# Quantifying and Enhancing Multi-modal Robustness with Modality Preference

### v1

@@ Abstract
Multi-modal models have shown a promising capability to effectively integrate information from various sources, yet meanwhile, they are found vulnerable to pervasive perturbations, such as uni-modal attacks and missing conditions. To counter these perturbations, robust multi-modal representations are highly expected, which are positioned well away from the discriminative multi-modal decision boundary. In this paper, different from conventional empirical studies, we focus on a commonly used joint multi-modal framework and theoretically discover that larger uni-modal representation margins and more reliable integration for modalities are essential components for achieving higher robustness. This discovery can further explain the limitation of multi-modal robustness and the phenomenon that multi-modal models are often vulnerable to attacks on the specific modality. Moreover, our analysis reveals how the widespread issue, that the model has different preferences for modalities, limits the multi-modal robustness by influencing the essential components and could lead to attacks on the specific modality highly effective. Inspired by our theoretical finding, we introduce a training procedure called Certifiable Robust Multi-modal Training (CRMT), which can alleviate this influence from modality preference and explicitly regulate essential components to significantly improve robustness in a certifiable manner. Our method demonstrates substantial improvements in performance and robustness compared with existing methods. Furthermore, our training procedure can be easily extended to enhance other robust training strategies, highlighting its credibility and flexibility.
@@ 1 Introduction
As data are often presented from different perspectives, like text, images, and audio, how to effectively exploit and integrate information from multiple sources becomes important. This has given rise to the concept of multi-modal learning, which serves as a potent approach that enables a more comprehensive understanding of complex concepts and facilitates more effective knowledge acquisition for different sources of information. Nowadays, multi-modal learning has demonstrated its remarkable ability in various tasks, such as video understanding (Yang et al., 2022), visual question answering (Antol et al., 2015), and emotion recognition (Tripathi et al., 2018).
@@
To address this, previous methods improve the training strategies to obtain a robust multi-modal model (Liang et al., 2021; Ding et al., 2021). Specifically, certain studies extend uni-modal robust training strategies, like adversarial training and mixup, to learn a discriminative multi-modal decision boundary (Li et al., 2022; Maheshwari et al., 2023). Others take steps like cross-modal alignment (Tian & Xu, 2021) to enhance the connection among modalities and obtain compact and robust multi-modal representation. However, even if they empirically improve robustness to some degree, they still lack in-depth theoretical analysis to understand the resilience of multi-modal models to perturbations, which is a vital aspect for safety-critical applications. More importantly, a universal phenomenon can be observed for these robust training methods that the adversarial attack on specific modalities could be more effective than others. As shown in Figure 1, the $\ell_{2}$-PGD attack with a range of radius is more effective on modality #$a$ than modality #$v$ for both Joint Training and three widely used robust training methods on the Kinetics Sounds dataset.
@@
Since the essential components of robustness, margins and integration, have an interrelation with each other, directly applying regulation is hard to guarantee higher certified robustness. To address this, we employ an orthogonal-based framework that formulates an alternative bound, which eliminates the interrelation and explicitly presents the integration. Building upon our theoretical analysis, we introduce a two-step Certifiable Robust Multi-modal Training (CRMT) procedure to ensure progressively superior robustness. Initially, we redefine uni-modal margins related to the reliability of the modality in this framework. Then we propose to regulate the unreliable modality by enlarging its margin, which can alleviate the imbalanced problem brought by modality preference. Further, our approach adjusts the integration of modalities considering the improvement of certified bound. These steps not only mitigate the large gap between robustness against attack on each modality but also credibly guarantee higher multi-modal robustness. To validate our method, we conduct extensive experiments to present the advanced robustness against both uni-modal and multi-modal attacks. Our contributions are as follows:
@@
    Through further analysis, we present evidence highlighting how multi-modal preference limits the multi-modal robustness and contributes to the vulnerability of multi-modal models towards specific modalities.
@@ 2 Related work
Recent studies have highlighted the vulnerability of deep neural networks (DNNs) to attacks and perturbations (Goodfellow et al., 2014; Madry et al., 2017), raising significant concerns regarding their deployment. With the presence of multiple modalities, the forms of perturbations include uni-modal attacks, multi-modal attacks (Schlarmann & Hein, 2023), and modality missing (Lee et al., 2023). One possible way to enable the models to resist these perturbations is to design robust training strategies to obtain a reliable decision boundary that is distanced from samples (Li et al., 2021). Certain methods, such as multi-modal adversarial training (Li et al., 2022) and incorporating uni-modal tasks (Ma et al., 2022), have proven pertinent in this context. On the other hand, studies focus on the latent consistency among modalities and propose to enhance modality interaction like alignment (Tsai et al., 2018) to obtain compact and robust representation (Bednarek et al., 2020). Among these empirical works, we identify a universal phenomenon that multi-modal models are commonly vulnerable to a certain modality (Liang et al., 2021), which currently lacks a comprehensive theoretical explanation. This motivates us to theoretically figure out the essential components determining the multi-modal robustness and explain the vulnerable modality phenomenon.
@@
Multi-modal learning is a significant approach for achieving a richer understanding across various tasks (Baltrušaitis et al., 2018; Jiang et al., 2021). However, even with the potential for a more comprehensive latent representation from multi-modal models (Huang et al., 2021), their performance might not always surpass that of the best uni-modal counterparts (Wang et al., 2020; Huang et al., 2022). Due to the inherent differences between modalities, a multi-modal model might prefer or lean towards a modality that is easier to learn, potentially overlooking others (Wu et al., 2022). In response, various strategies are proposed to strengthen the learning of uni-modalities, thereby enhancing the generalization of multi-modal model (Wang et al., 2020; Peng et al., 2022; Fan et al., 2023). However, these researches do not elucidate how this modality preference problem impacts robustness against perturbations, which is our main focus.
@@ 3.1 Preliminaries
In this section, we present the formulation of a multi-modal model considering a case with two modalities. Specifically, we consider a general $K$-way classification problem with a vectorized input sample $\bm{x}=\begin{bmatrix}\bm{x}^{(1)}\\ \bm{x}^{(2)}\end{bmatrix}\in\mathbb{R}^{d_{1}+d_{2}}$ consisting of two modalities, and a ground truth label $y\in[K]$. Our analyses focus on the robustness related to a certain sample, which can be further extended to encompass all $N$ samples across the dataset.
We consider the commonly used framework in multi-modal learning, the joint learning framework, where all of the modalities are jointly projected into the shared space for downstream tasks (Huang et al., 2022). Concretely, the encoder $\phi^{(m)}$ extracts the $m$-th modality’s input into representation. These uni-modal representations are concatenated to form a joint representation, which is then mapped to the output space using a linear classifier, where $W\in\mathbb{R}^{K\times(\mathrm{dim}(\phi^{(1)})+\mathrm{dim}(\phi^{(2)}))}$ and $\bm{b}\in\mathbb{R}^{K}$ are the weight matrix and bias respectively, and $\mathrm{dim}(\cdot)$ represents the dimension. The logits output of the multi-modal model can be denoted as $h(\bm{x})=W[\phi^{(1)}(\bm{x}^{(1)});\phi^{(2)}(\bm{x}^{(2)})]+\bm{b}=W^{(1)}\phi^{(1)}(\bm{x}^{(1)})+W^{(2)}\phi^{(2)}(\bm{x}^{(2)})+\bm{b}$, where $W^{(m)}\in\mathbb{R}^{K\times\mathrm{dim}(\phi^{(m)})}$ is the part of classifier $W$ related to the $m$-th modality.
@@ 3.2 Certified robustness for Multi-modal Model
Given an input $\bm{x}=\begin{bmatrix}\bm{x}^{(1)}\\ \bm{x}^{(2)}\end{bmatrix}$ with ground-truth label $y\in[K]$ and the closest label $j\neq y$, $\zeta_{j}^{(m)}(\bm{x}^{(m)})$ as the representation margin for $m$-th modality with Lipschitz constraint $\tau_{j}^{(m)}$, and the integration factor $c^{(m)}_{j}$. Define $\bm{x}^{\prime}=\begin{bmatrix}\bm{x}^{\prime(1)}\\ \bm{x}^{\prime(2)}\end{bmatrix}$ as the perturbed sample, and $\bm{x}-\bm{x}^{\prime}$ as the perturbation. The lower bound for the perturbation radius can be described as:
@@
The detailed proof can be found in the Appendix. Based on the certified bound above, we deduce that the multi-modal robustness depends on three primary factors: uni-modal representation margins $\zeta_{j}^{(m)}$, integration $c_{j}^{(m)}$, and bias difference $\beta_{j}$. Firstly, robustness increases proportionally with the enhancement of the uni-modal representation margin $\zeta_{j}^{(m)}$. Secondly, the integration $c_{j}^{(m)}$ is related to both uni-modal margins and the uni-modal Lipschitz constant, the expected integration for robustness requires considering both. Thus, a reasonable choice of integration factor can benefit higher robustness. Thirdly, when the sample is perturbed, the bias difference term $\beta_{j}$ stays invariant, since it depends on class $y$ and $j$ rather than the specific sample. Based on this fact, the bias difference is not considered in our analysis about model robustness. In a nutshell, we recognize that uni-modal margins and the integration of modalities are two essential components. In the following section, we will analyze how these essential components vary and influence the robustness of multi-modal models, especially under modality preference.
@@ 3.4 Certifiable Robust Multi-modal Training
Based on the above analysis, we target to improve the uni-modal representation margin $\zeta_{j}^{(m)}$ and adjust the integration of modalities $c_{j}^{(m)}$ as regulation for higher certified robustness. However, these regulations are intricately linked with the linear classifier $W$, potentially leading to conflicts in optimization objectives. As a result, stably enhancing the certified robustness presents challenges. Additionally, determining the Lipschitz constant for the margin between labels is computationally demanding. To address these challenges, we propose to adopt orthogonality within uni-modal linear classifiers as our framework. Detailedly, we ensure that in each modality, the class-specific vectors $\tilde{W}^{(m)}_{k\cdot},k\in[K]$ are unit and orthogonal. Since the valuable information between modalities is different, we apply the weight $\bm{a}^{(m)}\in\mathbb{R}^{K}$ to lead the model focusing on the more reliable modalities. Integrating these weights can enhance the model’s ability to effectively utilize the most valuable information from each modality. Therefore, learning of uni-modal representations and integration of modalities can be decoupled. The score corresponding to the $k$-th class can be expressed as:
@@
Given an input $\bm{x}=\begin{bmatrix}\bm{x}^{(1)}\\ \bm{x}^{(2)}\end{bmatrix}$ with ground-truth label $y\in[K]$ and the closest label $j\neq y$, the orthogonal classifier $\tilde{W}^{(m)}$, the modality-specific weight $\bm{a}^{(m)}$, the Lipschitz constant $\tilde{\tau}_{j}^{(m)}$, and the difference of the bias $\tilde{\beta}_{j}=\tilde{b}_{y}-\tilde{b}_{j}$. The lower bound for the perturbation radius with the orthogonal-based framework can be described as:
@@
The proof positions in the Appendix. With this lower bound, we can discuss designing a regulation method to explicitly enhance the certified robustness. Firstly, as analyzed in Section 3.3, the imbalance problem brought by modality preference impacts the unreliable modal representation margin. To explicitly regulate the uni-modal encoder and classifier independent from integration, we redefine the margin as the difference in uni-modal score between ground truth and the other label, which can also reflect the reliability of uni-modality. Then we propose to enlarge the relatively small margin through regularization, which is expressed as follows:
@@
which is detailed in the Appendix. Furthermore, we can adjust the integration of different modalities by enlarging the lower bound in Equation 10. Subsequently, we propose a two-step training procedure called Certifiable Robust Multi-modal Training (CRMT), which can credibly obtain a robust multi-modal model. The training procedure of CRMT is as follows:
@@
As shown in the Appendix, we further demonstrate that only one iteration of our method can achieve considerable robustness, without a huge consumption of training cost.
@@ 4.1 Setups
To make a thorough comparison, we evaluate our method on different datasets including Kinetics-Sounds (Audio + Vision) (Arandjelovic & Zisserman, 2017), UCF101 (Optical flow + RGB)  (Soomro et al., 2012), and VGGSound (Audio + Vision)  (Chen et al., 2020). We use the widely used backbone ResNet18 (He et al., 2016) as the uni-modal encoder, for each uni-modality. Then, the uni-modal representation is concatenated to form the multi-modal joint representation. Details about these datasets are presented in the Appendix.
@@
In this study, the comparison methods are selected to improve multi-modal learning and be suitable for the multi-modal joint training strategy. Comparison methods can be divided into two distinct groups: Firstly, methods address the imbalance problem caused by modality preference: Gradient Blending (GB) (Wang et al., 2020), On-the-fly Gradient Modulation with generalization enhancement (OGM) (Peng et al., 2022), Prototypical Modal Rebalance (PMR) (Fan et al., 2023). Secondly, methods aim at improving multi-modal robustness: Multi-Modal Adversarial Training (MMAT) (Li et al., 2022), Multi-modal mixup (Mixup) (Madry et al., 2017; Li et al., 2022), MinSim+ExFMem (MSEFM) (Tian & Xu, 2021). Our method can be extended to different training strategies, denoted as Certifiable Robust Multi-modal Training with Joint Training (CRMT-JT), CRMT with Adversarial Training (CRMT-AT), and CRMT with Mixup (CRMT-Mix).

### v2

@@ Abstract
Multi-modal models have shown a promising capability to effectively integrate information from various sources, yet meanwhile, they are found vulnerable to pervasive perturbations, such as uni-modal attacks and missing conditions. To counter these perturbations, robust multi-modal representations are highly expected, which are positioned well away from the discriminative multi-modal decision boundary. In this paper, different from conventional empirical studies, we focus on a commonly used joint multi-modal framework and theoretically discover that larger uni-modal representation margins and more reliable integration for modalities are essential components for achieving higher robustness. This discovery can further explain the limitation of multi-modal robustness and the phenomenon that multi-modal models are often vulnerable to attacks on the specific modality. Moreover, our analysis reveals how the widespread issue, that the model has different preferences for modalities, limits the multi-modal robustness by influencing the essential components and could lead to attacks on the specific modality highly effective. Inspired by our theoretical finding, we introduce a training procedure called Certifiable Robust Multi-modal Training (CRMT), which can alleviate this influence from modality preference and explicitly regulate essential components to significantly improve robustness in a certifiable manner. Our method demonstrates substantial improvements in performance and robustness compared with existing methods. Furthermore, our training procedure can be easily extended to enhance other robust training strategies, highlighting its credibility and flexibility. The code is available at https://github.com/GeWu-Lab/Certifiable-Robust-Multi-modal-Training.
@@ 1 Introduction
As data are often presented from different perspectives, like text, images, and audio, how to effectively exploit and integrate information from multiple sources becomes important. This has given rise to the concept of multi-modal learning, which serves as a potent approach that enables a more comprehensive understanding of complex concepts and facilitates more effective knowledge acquisition for different sources of information (Wei et al., 2022). Nowadays, multi-modal learning has demonstrated its remarkable ability in various tasks, including scene understanding (Antol et al., 2015; Yang et al., 2022; Li et al., 2022a), and emotion recognition (Tripathi et al., 2018; Chudasama et al., 2022).
@@
To address this, previous methods improve the training strategies to obtain a robust multi-modal model (Liang et al., 2021; Ding et al., 2021). Specifically, certain studies extend uni-modal robust training strategies, like adversarial training and mixup, to learn a discriminative multi-modal decision boundary (Li et al., 2022b; Maheshwari et al., 2023). Others take steps like cross-modal alignment (Tian & Xu, 2021) to enhance the connection among modalities and obtain compact and robust multi-modal representation. However, even if they empirically improve robustness to some degree, they still lack in-depth theoretical analysis to understand the resilience of multi-modal models to perturbations, which is vital for safety-critical applications. More importantly, a universal phenomenon can be observed for these robust training methods that the adversarial attack on specific modalities could be more effective than others. As shown in Figure 1, the $\ell_{2}$-PGD attack is more effective on modality #$a$ than modality #$v$ for both Joint Training and three widely used robust training methods on the Kinetics Sounds dataset.
@@
Since the essential components of robustness have an interrelation with each other, directly applying regulation is hard to guarantee higher robustness. To address this, we employ an orthogonal-based framework that formulates an alternative bound, which eliminates the interrelation and explicitly presents the integration. Building upon our theoretical analysis, we introduce a two-step Certifiable Robust Multi-modal Training (CRMT) procedure to ensure progressively superior robustness. Initially, we redefine uni-modal margins related to the reliability of the modality in this framework. Then we propose to regulate the unreliable modality by enlarging its margin, which can alleviate the imbalanced problem brought by modality preference. Further, our approach adjusts the integration of modalities considering the improvement of certified bound. These steps not only mitigate the large gap between robustness against attack on each modality but also credibly guarantee higher multi-modal robustness. To validate our method, we conduct extensive experiments to present the advanced robustness against both uni-modal and multi-modal attacks. Our contributions are as follows:
@@
    We present analyses highlighting how multi-modal preference limits the multi-modal robustness and contributes to the vulnerability of multi-modal models towards specific modalities.
@@ 2 Related work
Recent studies have highlighted the vulnerability of deep neural networks (DNNs) to attacks and perturbations (Goodfellow et al., 2014; Madry et al., 2017), raising significant concerns regarding their deployment. With the presence of multiple modalities, the forms of perturbations include uni-modal attacks, multi-modal attacks (Schlarmann & Hein, 2023), and modality missing (Lee et al., 2023). One possible way to enable the models to resist these perturbations is to design robust training strategies to obtain a reliable decision boundary that is distanced from samples (Li et al., 2021). Certain methods, such as multi-modal adversarial training (Li et al., 2022b) and incorporating uni-modal tasks (Ma et al., 2022), have proven pertinent in this context. On the other hand, studies focus on the latent consistency among modalities and propose to enhance modality interaction like alignment (Tsai et al., 2018) to obtain compact and robust representation (Bednarek et al., 2020). Among these empirical works, we identify a universal phenomenon that multi-modal models are commonly vulnerable to a certain modality (Liang et al., 2021), which currently lacks a comprehensive theoretical explanation. This motivates us to theoretically figure out the essential components determining the multi-modal robustness and explain the vulnerable modality phenomenon.
@@
Multi-modal learning is a significant approach for achieving a richer understanding across various tasks (Baltrušaitis et al., 2018; Jiang et al., 2021). However, even with the potential for a more comprehensive latent representation from multi-modal models (Huang et al., 2021), their performance might not always surpass that of the best uni-modal counterparts (Wang et al., 2020; Huang et al., 2022). Due to the inherent differences between modalities, a multi-modal model might prefer or lean towards a modality that is easier to learn, potentially overlooking others (Wu et al., 2022). In response, various strategies are proposed to strengthen the learning of uni-modalities, thereby enhancing the generalization of multi-modal model (Wang et al., 2020; Peng et al., 2022; Fan et al., 2023; Xu et al., 2023b). However, these researches do not elucidate how this modality preference problem impacts robustness against perturbations, which is our main focus.
@@ 3.1 Preliminaries
We consider a general $K$-way classification problem with a vectorized input sample $\bm{x}=(\bm{x}^{(1)},\bm{x}^{(2)})\in\mathbb{R}^{d_{1}+d_{2}}$ consisting of two modalities, and a ground truth label $y\in[K]$. We consider the commonly used joint learning framework in multi-modal learning, where all of the modalities are projected into the shared space for downstream tasks (Huang et al., 2022). Concretely, the uni-modal representations are extracted by encoders $\phi^{(m)}$ and concatenated to form the joint representation, which is then mapped to the output space using a linear classifier, where $W\in\mathbb{R}^{K\times(\mathrm{dim}(\phi^{(1)})+\mathrm{dim}(\phi^{(2)}))}$ and $\bm{b}\in\mathbb{R}^{K}$ are the weight matrix and bias respectively, and $\mathrm{dim}(\cdot)$ represents the dimension. The logits output of the multi-modal model can be denoted as $h(\bm{x})=W[\phi^{(1)}(\bm{x}^{(1)});\phi^{(2)}(\bm{x}^{(2)})]+\bm{b}=W^{(1)}\phi^{(1)}(\bm{x}^{(1)})+W^{(2)}\phi^{(2)}(\bm{x}^{(2)})+\bm{b}$, where $W^{(m)}\in\mathbb{R}^{K\times\mathrm{dim}(\phi^{(m)})}$ is the part of classifier $W$ related to the $m$-th modality.
@@ 3.2 Certified robustness for Multi-modal Model
Given an input $\bm{x}$ with ground-truth label $y\in[K]$ and the closest label $j\neq y$, $\zeta_{j}^{(m)}(\bm{x}^{(m)})$ as the representation margin for $m$-th modality with Lipschitz constraint $\tau_{j}^{(m)}$, and the integration factor $c^{(m)}_{j}$. Define $\bm{x}^{\prime}$ as the perturbed sample, and $\bm{x}-\bm{x}^{\prime}$ as the perturbation. The lower bound for the perturbation radius can be described as:
@@
The detailed proof can be found in the Appendix 7.1. Based on the certified bound above, we deduce that the multi-modal robustness depends on three primary factors: uni-modal representation margins $\zeta_{j}^{(m)}$, integration $c_{j}^{(m)}$, and bias difference $\beta_{j}$. Firstly, robustness increases proportionally with the enhancement of the uni-modal representation margin $\zeta_{j}^{(m)}$. Secondly, the integration $c_{j}^{(m)}$ is related to both uni-modal margins and the uni-modal Lipschitz constant, the expected integration for robustness requires considering both. Thus, a reasonable choice of integration factor can benefit higher robustness. Thirdly, when the sample is perturbed, the bias difference term $\beta_{j}$ stays invariant, since it depends on class $y$ and $j$ rather than the specific sample. Based on this fact, the bias difference is not considered in our analysis of model robustness. In a nutshell, we recognize that uni-modal margins and the integration of modalities are two essential components. In the following section, we will analyze how these essential components vary and influence the robustness of multi-modal models, especially under modality preference.
@@ 3.4 Certifiable Robust Multi-modal Training
Based on the above analysis, we target to improve the uni-modal representation margin $\zeta_{j}^{(m)}$ and adjust the integration of modalities $c_{j}^{(m)}$ as regulation for higher certified robustness. However, these regulations are intricately linked with the linear classifier $W$, potentially leading to conflicts in optimization objectives. As a result, stably enhancing the certified robustness presents challenges. Additionally, determining the Lipschitz constant for the margin is computationally demanding. To address these challenges, we propose to adopt orthogonality  (Huang et al., 2018) within uni-modal linear classifiers as our framework. Detailedly, we ensure that in each modality, the class-specific vectors $\tilde{W}^{(m)}_{k\cdot},k\in[K]$ are unit and orthogonal. Since the valuable information between modalities is different, we apply the weight $\bm{a}^{(m)}\in\mathbb{R}^{K}$ to lead the model focusing on more reliable modalities. Integrating these weights can enhance the model’s ability to effectively utilize the valuable information from each modality. Therefore, learning of uni-modal representations and integration of modalities can be decoupled. The score corresponding to the $k$-th class can be expressed as:
@@
Given an input $\bm{x}$ with ground-truth label $y\in[K]$ and the closest label $j\neq y$, the orthogonal classifier $\tilde{W}^{(m)}$, the modality-specific weight $\bm{a}^{(m)}$, the Lipschitz constant $\tilde{\tau}_{j}^{(m)}$, and the difference of the bias $\tilde{\beta}_{j}=\tilde{b}_{y}-\tilde{b}_{j}$. The lower bound for the perturbation radius with the orthogonal-based framework can be described as:
@@
See Appendix 7.2 for proof. With this lower bound, we can discuss designing a regulation method to explicitly enhance the certified robustness. Firstly, as analyzed in Section 3.3, the imbalance problem brought by modality preference impacts the unreliable modal representation margin. To explicitly regulate the uni-modal encoder and classifier independent from integration, we redefine the margin as the difference in uni-modal score between ground truth and the other label, which can also reflect the reliability of uni-modality. Then we propose to enlarge the relatively small margin through regularization, which is expressed as follows:
@@
which is detailed in the Appendix 7.3. Furthermore, we can adjust the integration of different modalities by enlarging the lower bound in Equation 10. Subsequently, we propose a two-step training procedure called Certifiable Robust Multi-modal Training (CRMT), which can credibly obtain a robust multi-modal model. The training procedure of CRMT is as follows:
@@
As shown in Appendix 8.3, we further demonstrate that only one iteration of our method can achieve considerable robustness, without a huge consumption of training cost.
@@ 4.1 Setups
We evaluate our method on different datasets including Kinetics-Sounds (Audio + Vision) (Arandjelovic & Zisserman, 2017), UCF101 (Optical flow + RGB)  (Soomro et al., 2012), and VGGSound (Audio + Vision)  (Chen et al., 2020). We use the backbone ResNet18 (He et al., 2016) as the encoder for each uni-modality. Details about these datasets are presented in Appendix 8.1.
@@
In this study, the comparison methods are selected to improve multi-modal learning and be suitable for the multi-modal joint training strategy. Comparison methods can be divided into two distinct groups: Firstly, methods address the imbalance problem caused by modality preference: Gradient Blending (GB) (Wang et al., 2020), On-the-fly Gradient Modulation with generalization enhancement (OGM) (Peng et al., 2022), Prototypical Modal Rebalance (PMR) (Fan et al., 2023). Secondly, methods aim at improving multi-modal robustness: Multi-Modal Adversarial Training (MMAT) (Li et al., 2022b), Multi-modal mixup (Mixup) (Madry et al., 2017; Li et al., 2022b), MinSim+ExFMem (MSEFM) (Tian & Xu, 2021). Our method can be extended to different training strategies, denoted as Certifiable Robust Multi-modal Training with Joint Training (CRMT-JT), CRMT with Adversarial Training (CRMT-AT), and CRMT with Mixup (CRMT-Mix).